### v1

# Data-Model Co-Evolution: Growing Test Sets to Refine LLM Behavior

Abstract.
A long-standing challenge in machine learning has been the rigid separation between data work and model refinement, enforced by slow fine-tuning cycles. The rise of Large Language Models (LLMs) overcomes this historical barrier, allowing applications developers to instantly govern model behavior by editing prompt instructions. This shift enables a new paradigm: data-model co-evolution, where a living test set and a model’s instructions evolve in tandem. We operationalize this paradigm in an interactive system designed to address the critical challenge of encoding subtle, domain-specific policies into prompt instructions. The system’s structured workflow guides people to discover edge cases, articulate rationales for desired behavior, and iteratively evaluate instruction revisions against a growing test set. A user study shows our workflow helps participants refine instructions systematically and specify ambiguous policies more concretely. This work points toward more robust and responsible LLM applications through human-in-the-loop development aligned with local preferences and policies.
Large language models, test set, model behavior, human-AI interaction, human-in-the-loop workflows, responsible AI
$c$s: Human-centered computing Interactive systems and toolsccs: Computing methodologies Natural language generation
Figure 1. Overview of the data-model co-evolution paradigm. Traditionally, prompt instructions (model) and test set examples (data) have been developed separately with slower model tuning cycles. With rapid iteration, however, these two artifacts can now evolve together. Our approach operationalizes this opportunity by treating each discovered failure as a test case and validating each prompt instruction revision against the growing test set, enabling systematic refinement.
Content warning: This paper includes some text examples containing offensive language.
1. Introduction
Large Language Models (LLMs) are driving a new wave of application development. This empowers people, including those without programming expertise, to direct model behavior through natural language instructions, a practice known as prompt engineering (Zamfirescu-Pereira et al., 2023; Subramonyam et al., 2025). However, it comes with a critical challenge: a general-purpose model inherently lacks the local context and nuanced policies of a specific application (Palla et al., 2025; In et al., 2025). For instance, a small online community may wish to foster playful banter by allowing specific slang, a nuance lost on a strict, one-size-fits-all content moderation filter from a major tech vendor, such as OpenAI and Google. This gap forces people into a reactive cycle of prompt tinkering, making small, ad-hoc edits and spot-checking them against a few examples to patch failures at the margins.
This ad-hoc approach reflects a deeper mismatch between the new speed of development and the old tools at hand. Many current practices for evaluation, such as building static test sets, are holdovers from the traditional machine learning era (Sambasivan et al., 2021). Back then, the slow, costly process of model fine-tuning necessitated a strict separation between data curation and model refinement. In today’s interactive landscape, where application developers can iterate on prompts in seconds, this inherited separation has become a major source of friction, preventing a truly fluid and responsive development process.
This friction is compounded by a new bottleneck: the sheer difficulty of authoring comprehensive test data for the long tail of modern LLM-based applications. Manually anticipating the diverse ways users will interact with generative models is a formidable task. To compensate, specialized and labor-intensive efforts have emerged. Red teaming, for example, uncovers critical vulnerabilities but often operates as a siloed, episodic exercise whose findings rarely integrate back into a persistent test suite (Feffer et al., 2024; Kiela et al., 2021; Quaye et al., 2024). Similarly, policy specifications or “constitutions” provide high-level guidance but tend to be static artifacts, difficult to update against a continuous stream of new edge cases (Bai et al., 2022; Petridis et al., 2024). Whether through ad-hoc tinkering, episodic testing, or static policies, a crucial feedback loop between discovering failures and refining behavior remains broken. We argue that the problem is the separation of data and model, where prompting is fast, but evaluation is slow.
To address this, we introduce data–model co-evolution, a new paradigm where the instructions (the model) and the test examples (the data) are treated as intertwined artifacts that grow together. Crucially, we define the ‘model’ from the application developer’s perspective: it is the editable specification, typically a system prompt that encodes policies and instructions that the part of the system they can actually edit to control its behavior. In our approach, this model helps discover its own blind spots, which are then immediately transformed into persistent test cases. Each revision is then tested against this growing set of real-world examples, creating the tight feedback loop that current practices lack.
To operationalize this paradigm, we present an interactive system that supports users to refine prompt instructions and create test sets together for generative applications. Users begin with an initial policy and prompt. The system proposes candidate user inputs, including those likely to be ambiguous under the given instruction, and runs the target model to reveal behavior. When the output is incorrect, the system engages the user in a dialog: they mark the example and articulate a short rationale. To test generalization to other examples, the system generates small neighborhoods of similar cases for quick labeling. These labeled slices both grow the test set and inform revisions to the prompt. Users are then directed to compare model responses before and after revision, with an LLM-as-judge (Zheng et al., 2023) seeded by the user’s own labels. Crucially, all labeled data persist, yielding a living, incrementally expanding test set.
Our approach is grounded in formative interviews with five practitioners, which revealed a clear need for workflows that better integrate systematic, data-driven evaluation into the development loop. We then evaluated our system in a controlled study (N=12) against a baseline. Our workflow led participants to produce more detailed instructions with explicit boundaries and examples, build larger and richer test sets rapidly and systematically, induced more consistent changes in model behavior, and reported higher satisfaction compared to a baseline.
Our contributions are as follows:
  • A new paradigm of data–model co-evolution, positioning test-set growth and specification refinement as incrementally co-evolving artifacts in LLM application development, informed by formative interviews with practitioners.
  • An interactive system that embodies this paradigm through structured workflows that support edge case discovery, rationale elicitation, neighborhood probing, and regression testing against a living test set.
  • Empirical insights from a controlled lab study (N=12), demonstrating that our workflow produces more specific prompt instructions, larger test sets, more consistent changes to model behavior, and higher user satisfaction than a baseline.
Our work contributes to HCI and human-AI interaction by advancing methods for democratizing LLM development and enabling more responsible, policy-aware use of AI. By shifting from prompt tinkering or episodic red teaming toward continuous, human-in-the-loop test-set growth, we show how interactive systems can help diverse developers—not only AI experts—make sense of subtle application-specific norms, codify them incrementally, and build more robust and accountable LLM applications.
2. Related Work
We connect our work to five strands: human-centered prompt design, interactive machine learning and labeling, data iteration and test-driven evaluation, bi-directional human-AI alignment, and human-in-the-loop red teaming.
2.1. Human-Centered Prompt and Policy Design
Recent HCI research on LLMs and prompt engineering has framed prompts as design artifacts that encode requirements and governance choices. For example, ConstitutionMaker (Petridis et al., 2024) demonstrates how user critiques can be transformed into enforceable principles, while ROPE (Ma et al., 2025) highlights the role of prompts as requirement specifications. Policy Maps (Lam et al., 2024) treats policy design as a cartographic practice, providing interactive maps that help practitioners navigate and demarcate regions of acceptable and unacceptable LLM behaviors. PromptHive (Reza et al., 2025) illustrates how domain experts can co-create prompts. Together, these projects emphasize that prompts are sites where human values and policies are concretized.
Complementing these policy-oriented perspectives, other work has built interactive systems to help practitioners explore, compare, and refine prompts. PromptIDE (Strobelt et al., 2022) provides visual analytics support for prompt iteration, EvalLM (Kim et al., 2024) introduces workflows for comparing different prompt versions through structured evaluation, and ChainForge (Arawjo et al., 2024) enables rapid experimentation with LLM prompt chains. In addition, researchers have explored alternative familiar form factors, spreadsheets, for supporting prompt iteration in the context of data labeling and classification (He et al., 2025; Qian et al., 2025).
In parallel, the Natural Language Processing (NLP) community has largely pursued automatic prompt optimization, treating prompt design as a black-box search problem (Cheng et al., 2024). Recent techniques include black-box optimization for alignment, multi-branch optimization strategies (Yang et al., 2024), and embedding-based prompt search (Cheng et al., 2023). These methods improve performance but rarely help stakeholders articulate policies or test boundary cases.
2.2. Interactive Machine Learning and Labeling
Interactive machine learning (IML) has long shown that humans and models can co-adapt through iterative data curation (Amershi et al., 2014; Dudley and Kristensson, 2018). The Machine Teaching paradigm argued for exposing teaching levers, so non-experts can shape model behavior with principled, design-forward workflows. (Simard et al., 2017).
Recent mixed-initiative tools leverage LLMs for data annotation. PaTAT (Gebreegziabher, 2025) supports qualitative coding by synthesizing interpretable rules from a user’s codes, and Wang et al. (Wang et al., 2024) proposes a collaboration workflow where LLMs produce labels and explanations, a verifier assesses their quality, and human annotators re-annotate lower-quality ones.
Beyond labeling workflows, prior HCI research has shown that asking users to articulate why a model’s output is wrong can play a critical role in debugging and refinement. Kulesza et al. (Kulesza et al., 2015) used the term explanatory debugging for explanations that helps end users form more accurate mental models of how a system works, leading to more effective corrections. Later studies argued that explanations can support tasks beyond model understanding (Kim et al., 2023; Liao et al., 2020). Our workflow extends this tradition by eliciting rationales for failures, leading to prompt revisions and test set improvements.
In parallel, the data-centric AI movement argues that systematic data design and curation beyond model tuning drive performance, advocating practices for engineering data quality (Whang et al., 2023). Weak-supervision frameworks like Snorkel (Ratner et al., 2017) excel at throughput and coverage, but often miss opportunities for interactive elicitation.
2.3. Data Iteration and Test-Driven Development
A large body of HCI work argues that the importance of datasets in ML development. Chameleon (Hohman et al., 2020) shows how practitioners can attribute performance changes to data, comparing features, splits, and successive dataset versions through interactive visualizations. A series of interview studies further document the difficulty and iterative nature of data work in machine learning development in practice (Sambasivan et al., 2021; Holstein et al., 2019; Crisan and Fiore-Gartland, 2021; Qian et al., 2025).
Behavioral testing frameworks operationalize this stance by turning errors into executable tests. CheckList (Ribeiro et al., 2020) provides a tooling to ideate perturbations and generate broad suites of checks, showing that teams uncover substantially more actionable bugs than with ad-hoc evaluation. There also exist the use of a domain-specific language to define precise error groups and run counterfactual rewrites (Wu et al., 2019, 2020). Zeno (Cabrera et al., 2023a) develops an interactive web UI to slice data, specify behaviors, and track results across tasks. AdaTest (Ribeiro and Lundberg, 2022) takes human feedback to generate unit tests for debugging NLP models. We adopt the behavioral stance by creating slices from incorrectly generated examples, but center policy-boundary ambiguity with LLM generation, and let users compare new versions to older ones for assessment as well.
2.4. Bidirectional Human-AI Alignment and Criteria Drift
The notion of alignment has become central in AI research, referring to steering model behavior toward human goals through reinforcement learning from human feedback (RLHF) or similar algorithms (Christiano et al., 2017; Ouyang et al., 2022). HCI work, however, stresses that humans and AI must develop mutual understanding and continuously adjust to one another, a view captured in the idea of bidirectional alignment (Shen et al., 2025). This perspective includes studies on user-centered explanations for calibrating trust and mitigating overreliance (Vasconcelos et al., 2023; Kim et al., 2023), work that help develop users’ mental models about model behavior (Cabrera et al., 2023b, c), and investigations in high-stakes domains, where experts actively negotiate AI recommendations (Sivaraman et al., 2023).
Alongside this, alignment is tied to evolving criteria for evaluation. Building on work that highlights how evaluation frameworks must reflect user-defined criteria (Kim et al., 2024), Shankar et al. (Shankar et al., 2024) describe criteria drift, where users’ standards change as they interact with model outputs. Our work differs by focusing not only on criteria for judging static data outputs but also on refining the underlying policy specifications.
2.5. Human’s Roles in Red Teaming
Red teaming has emerged as a central strategy for stress-testing generative models. Automated pipelines exist for generating adversarial inputs (Perez et al., 2022), but researchers emphasize their limited creativity and coverage, making human input indispensable (Feffer et al., 2024). To scaffold human creativity at scale, the Adversarial Nibbler challenge (Quaye et al., 2024) recruited diverse communities to expose vulnerabilities in text-to-image generation, based on the Dynabench (Kiela et al., 2021) project. In addition, Yeh et al. (Yeh et al., 2025) and Deng et al. (Deng et al., 2025) leverages visual analytic solutions to help red-teaming professionals explore the space of adversarial prompts.
Our approach builds on this line of work but shifts the focus: rather than treating red teaming purely as dataset exploration, we emphasize iterative engagement with models through policy-grounded prompts. Prior red-teaming research often assumes a division of labor—dataset curators probe models while developers remain separate, but as LLM development becomes democratized, this separation may not hold. Domain experts in vertical applications, small companies, and individual developers increasingly need to red-team their own systems without the infrastructure of large organizations. Our workflow supports them by providing structured methods to capture community-specific rules and policies, extending red-teaming practices beyond benchmark collection toward a process accessible even to those without deep technical expertise.
3. Formative Study
To ground our work in real-world practices, we conducted a formative study to understand the challenges industry practitioners face when developing and refining LLM-based applications. We sought to answer: how do teams currently discover, diagnose, and address failures in their LLM systems, and how do these failures feed into their processes for improving their applications?
3.1. Participants and Procedure
We conducted 60-minute semi-structured interviews over video conferencing with five practitioners from various industries. The participants came from diverse professional roles, including data scientists, software engineers, product managers, and AI researchers with two to eleven years of experience in their fields. They were actively building and deploying a range of LLM applications, including natural-language-to-SQL query systems, call center dialogue summarization and classification, banking chatbots, and AI-assisted planning applications. Our interviews focused on their day-to-day workflows for development, evaluation, and iteration. The interviews were recorded, transcribed, and coded through a thematic analysis. The study was approved by our institution’s IRB.
3.2. Findings
We identified four key findings that highlight a gap between current practices and the need for more systematic tooling.
3.2.1. Development is a Process of Constant, Unending Iteration
The most consistent theme across all participants was that building a reliable LLM application is not a one-time task but a continuous cycle of refinement. Practitioners described a reactive process where “unknown unknowns” constantly emerge. Early deployments often resembled extended beta tests, with colleagues, testers, or even early users surfacing unexpected cases that had never been anticipated during initial design. Each new discovery was then folded back into the workflow, typically by editing the prompt instruction with a new policy rule, followed by quick evaluation to see whether the change generalized.
“Once people actually used it, we kept getting “this doesn’t work” cases. We took those and fed them back into the model, and kept iterating in small steps until the final test.” (P1)
“For two months, all I did was tweak and tune prompts. Accuracy went up, but at some point we just had to settle.” (P2)
This cycle of iteration–collecting new cases, updating prompts, and testing again–was described as never-ending. Below, we take a closer look at how practitioners actually conduct this iterative process.
3.2.2. Prompt Engineering is the Preferred Modality for Rapid Iteration
Many practitioners expressed a preference for prompt engineering over fine-tuning as their primary development method. While some teams with significant resources used fine-tuning, the majority favored prompting due to its speed and cost.
“We handled everything through prompts […] if we could fix it with prompting, we did.” (P2)
“We considered embedding filters […], but due to latency we found it more effective to encode guardrails in the system prompt for fast response to new attacks.” (P3)
As model capabilities continue to improve, the challenge has shifted from enhancing raw performance to aligning outputs with nuanced, application-specific requirements. Practitioners noted that capability is often “good enough,” making it more important to encode policies, edge cases, and organizational constraints directly through prompts. In this sense, prompt engineering becomes not just a convenient shortcut but the central mechanism for cost-effective, policy-aligned iteration.
3.2.3. Test Sets are Diverse, Application-Specific, and Often Ad-Hoc
Practitioners test their applications against a wide variety of inputs, but these test sets are often curated informally. The content of these tests is highly dependent on the application’s context:
  • Difficulty Levels: Teams developing tasks such as natural language to SQL assistants often stratified examples into “easy,” “medium,” and “hard” to probe capability stepwise and to guard against regressions. Participants emphasized distributing tests across these bands to check that simple cases remain reliable, and truly difficult edge cases surface specification gaps.
  • Out-of-Scope Inputs: For customer-facing systems (e.g., banking chatbots), a recurring concern was how the model handles irrelevant or nonsensical queries. Participants noted that real users often “try anything” either out of curiosity about the AI or to probe its boundaries. Tests in this category therefore evaluate both detection (is the request off-domain?) and response style (is the refusal informative and consistent).
  • Problematic and Sensitive Inputs: For enterprise deployments, participants reported dedicating significant effort to guardrail tests targeting harmful, biased, or brand-damaging outputs. Beyond generic safety, several vendors had client-specific red lines, for example, mentioning competitor products. Test sets here can cover examples that are often used in the context of red teaming, e.g., abusive or harassing language, and sexually explicit requests. Because the cost of failure is high, these tests were described as high priority, and when teams felt they lacked sufficient expertise, they sometimes relied on external vendors.
These collections of test cases were often informal artifacts, e.g., just a list of examples in a shared document, lacking a systematic structure for evaluation.
3.2.4. Evaluation Dichotomy Between Rapid Gut-Checks and Costly Metrics
We found that practitioners’ evaluation practices are split into two distinct modes: rapid, intuitive checks and formal, costly benchmarking.
For the frequent, small-scale iterations typical of prompt engineering, the dominant method is the “gut-check”—“In the end, what really mattered was more of a gut feeling—like, ‘this version feels more concrete and actually covers the things we need.”’ [P4]. Practitioners make a small change and then manually observe a handful of outputs to get a “feel” for the impact. This approach is valued for its speed. As one participant described, “Three developers ran one–two-hour jam sessions, throwing queries and checking logs. Roughly  100 queries in total; if results looked good, we locked in the system prompt.” [P3]. This informal, qualitative assessment is the engine of the day-to-day development loop.
However, to gain higher confidence or to justify changes to stakeholders, this intuitive sense is not enough. In these situations, teams turn to a more formal process: curating a test set, manually labeling each output, and calculating quantitative metrics. Providing the ”hard numbers” needed to demonstrate progress to clients or leadership.
This creates a central tension in their workflow. While practitioners desire the rigor of quantitative evaluation, its high cost in time and labor makes it impractical for every iteration. As one participant put it, “Human review by the representatives wasn’t feasible on every cycle. Practically, we could only do it about once a month.” [P4]. This dichotomy forces teams into a difficult compromise: they perform the majority of their iterations using fast but low-confidence gut-checks, punctuated by infrequent, high-effort formal evaluations. This leaves them without a continuous, clear signal of whether their iterative changes are truly improving the system.
3.3. Design Implications
Our formative study revealed a clear gap: while practitioners are locked in a rapid, iterative loop of prompt refinement, they lack systematic tools to support this process. These findings motivated the design of a workflow grounded in a new paradigm.
  • Systematize Iteration: The constant, reactive cycle practitioners described (Finding 1 & 2) requires a workflow that is both fast and structured. This calls for providing a tight, interactive loop for discovering a failure and immediately refining the specification.
  • Structure the Test Set: To address the ad-hoc nature of testing (Finding 3), the test set should be treated as a first-class citizen. It provides a formal structure for accumulating diverse test cases—including the edge cases discovered through generation—creating a durable, living benchmark.
  • Accelerate Evaluation: To alleviate the burden of manual review (Finding 4), evaluation can be accelerated with side-by-side comparisons between instruction versions (Kim et al., 2024; Zheng et al., 2023), while still allowing for human oversight. The automated accuracy metrics provide the quantitative feedback practitioners need for tracking progress and for persuasion.
Toward a New Paradigm. Synthesizing these implications, we identified a fundamental disconnect in current workflows: the prompt specification (model) and the test user inputs (data) are treated as separate, informally managed artifacts. A change to one does not systematically inform or validate the other. This insight led us to propose a new paradigm for LLM development: data–model co-evolution. In this paradigm, the test set and the specification are not independent but are treated as intertwined artifacts that must grow and evolve in tandem. Each discovered failure (data) must durably impact the specification (model), and each change to the specification must be validated against all accumulated data. Our system is a direct embodiment of this paradigm.
Figure 2. System interface for the Co-Evolution Workflow. The left panel supports authoring and tracking prompt instructions across versions. The right panel manages data: the top section maintains a living test set with model responses from both prior and current prompt instruction versions, while the bottom section displays generated candidate inputs for user review. Together, these panels embody the co-evolution principle by tightly coupling specification revisions with test set examples. The examples in this figure, originally written in Korean, were used in the user study and translated into English.
4. Interactive Co-Evolution Workflow
To operationalize our paradigm of data–model co-evolution, we designed and built an interactive system for refining LLM behavior. We first describe the design principles that guided the workflow (4.1), then introduce the interface (4.2), and finally explain the workflow in detail (4.3).
4.1. Design Principles
Before presenting the principles, we first clarify our terminology. In this work. In the context of LLM-based applications, we use the term model in a broadened sense. Specifically, the model refers to the prompt instruction (a.k.a., system prompt) that encodes the intended policy, which is concatenated with a given user input and sent to the LLM API to produce a response. By contrast, we use the term data to denote end users’ inputs which invoke the model to produce different outputs. In other words, the model remains stable as the governing specification, while the data varies with every user interaction. With this in place, we describe our key design principles:
4.1.1. From Model to Data: Using the Specification to Discover Its Own Blind Spots.
The refinement process begins with the current model (i.e., prompt instruction). Its weaknesses are revealed not in the average case, but in specific, unforeseen edge cases. We leverage the powerful data synthesis capabilities of modern LLMs to find these weaknesses. The system uses the current prompt instruction to generate challenging inputs, which surfaces potential failure modes and helps the user ideate and anticipate problematic cases.
4.1.2. From Data to Model: Grounding Specification Refinements in Concrete Evidence.
Once the edge cases (data) are surfaced, they become the foundation for refining the prompt instruction (model). This principle ensures that changes are not based on vague intuition but on concrete evidence. As a user confronts a generated failure, they are prompted to label it and articulate a rationale. This step can help clarify their own internal policy. It is then used to guide the update to the prompt instruction.
4.1.3. A Living Test Set:
Each discovered case is preserved in a growing test set that reflects the current state of the model. This evolving benchmark enables users to assess overall progress, detect regressions, and ensure systematic improvement.
4.1.4. LLM-Assisted, Human-Governed Workflow
Throughout the workflow, the LLM acts not only as the object of refinement but also as a collaborative assistant. It assists at every stage: suggesting edge cases, proposing rationales for failures, generating similar examples for generalization, and even suggesting prompt revisions. Yet all suggestions remain editable, keeping humans the final authority while accelerating their process.
4.1.5. Scaffolding Generalization: From a Single Case to a General Rule:
The workflow aims to support translating the insight from single failures into generalizable rules. Our workflow explicitly scaffolds this transition by allowing users to test that a rationale against similar cases, before committing it to the model update. This helps transform a reactive fix (”this one output was bad”) into a proactive policy enhancement (”this class of outputs is undesirable, and here is a rule to prevent it”), building users confidence and leading to more robust models.
4.2. System Interface
Figure 2 shows the interface. It consists of two main columns that physically separate the model specification from the data. The left panel is dedicated to editing and tracking the prompt instruction, while the right panel is for discovering and curating the test data.
4.2.1. The Prompt Instruction (Model) Panel
The left side of the interface is the control center for the prompt instruction.
Prompt Instruction Editor: At the top is a text editor where developers write and refine the instructions that guide the LLM’s behavior.
Version History: Below the editor, a history view lists all saved versions of the prompt instruction. It also serves as a performance dashboard. Each version is visualized with a percentage bar chart to show the performance metric of its overall accuracy calculated against the current test set. This allows users to see at a glance whether a revision improved performance or caused an issue. Users can easily revert to or compare against previous versions.
4.2.2. The Data Panel: Discovering and Accumulating Evidence
The right side of the interface is dedicated to the data that informs and validates the prompt specification. It consists of two tables.
The Test Set: The main component is the table representing the test set. It is a collection of labeled examples that serve as the ground truth for the application’s policy. Its columns display user inputs, model responses, Good/Bad ratings, and the corresponding rationales. When a new prompt version is evaluated, this table updates to show responses from both the old and new versions side-by-side, making the impact of any change immediately apparent.
Generated Examples: Below the test set, a smaller table serves as a staging area for examples synthesized by LLMs. This is where new, unlabeled edge cases are surfaced for the user to review. When an example from this area is labeled, it is formally promoted into the test set above.
4.3. Co-Evolution Workflow in Action
A user’s interaction with the system follows an iterative cycle designed to discover failures, generalize from them, and refine the model’s specification against the test set, which is illustrated in Figure 3. Below, we describe each step of the workflow, using the example of a banking customer-support chatbot that should only handle finance-related queries and politely redirect queries that are out of scope.
Figure 3. Overview of the Co-Evolution workflow: (1) the system synthesizes potential user inputs that are likely hard or ambiguous; (2) users examine these cases, label failures, and provide rationales as hypotheses for model errors; (3) neighborhood probing generates similar inputs to test whether the rationale generalizes; (4) the system proposes a revised prompt instruction grounded in the rationale and neighborhood labels; and (5) the expanded test set evaluates model responses across both prior and updated prompt instruction versions. This structured loop scaffolds the shift from a single error to a broader policy rule.
Step 1: Discovering Failures. The workflow begins in the Generated Examples panel, where the system generates candidate user inputs intended to be “hard” cases under the current prompt instruction. These examples are generated by an LLM with the explicit goal of probing boundaries that are likely ambiguous or underspecified. Users inspect the model’s responses to these inputs and labels them as Good or Bad. For example, the chatbot might receive the input “How should I manage my credit score while preparing for job applications?”. If the AI refuses to answer on the grounds that this is an “employment-related” question, the user may label this as Bad, reasoning that credit score management is central to their banking support even if the query also mentions job preparation. While some applications might choose to withhold an answer in such cases, our scenario reflects a policy preference to provide guidance as long as the query does not cross into problematic territory.
Step 2: Articulating a Rationale. When a user labels an example as Bad, the system prompts them to provide a rationale explaining why the behavior was incorrect. Instead of leaving this as a free-form burden, the interface provides two LLM-suggested options, but the users can also write their own. For example, they might enter “Questions about credit scores should be answered, even if they are mentioned in the context of employment or other life events.” This articulation transforms an implicit judgment into an explicit principle, shifting the mindset from reactive correction (“this answer is wrong”) to policy refinement (“here is the scope rule that the chatbot should follow”).
Figure 4. Neighborhood validation step in the workflow. The top section shows the original selected example, along with the user’s label and written rationale. Below, the system provides semantically similar examples that vary slightly from the original, enabling the user to test whether their rationale holds across a neighborhood of cases. As users label these examples, they actively reflect on the policy boundary, and the results directly inform the next step of revising the prompt instruction. The validated examples are then added to the test set, where they continue to be used in subsequent evaluations.
Step 3: Generalizing the Rationale with Neighborhood Probing. The aim of this step is to move from a single failure to a robust, reusable rule. Our workflow supports neighborhood probing: given the user’s rationale, the system synthesizes a small set of semantically similar inputs that preserve the same triggers while varying context. Clicking a button on the rationale popup opens a larger popup for this purpose. As shown in Figure 4, users can see and label these cases to see if the rationale consistently predicts undesired behavior. If the model mishandles all of them, the rationale is confirmed as broadly relevant; if not, the developer may refine the rationale to be more specific. These examples will be added to the test set. We crafted a prompt that generates three new examples as below:
Generate three new user inputs to verify if the failure described in the Human Rationale is reproducible,
without widening scope.
Context
- AI Instruction: {prompt_instruction}
- Original User Input: {user_input}
- AI Response: {ai_response}
- Human Rationale (root cause & semantic triggers): {human_rationale}
Precedence (for any conflict)
1) Human Rationale (triggers & scope)
2) Original inputs style anchors
3) Reference-only info (surface flavor only; ignore if it broadens scope)
Anchors to preserve
- Same semantic trigger(s) as in Human Rationale (paraphrase allowed only if meaning identical)
- Style anchors of the original: language/register, salient markers
Step 4: Revising the Specification. Once the neighborhood is labeled, the dialog proceeds to the next step: prompt instruction revision. The system synthesizes the original failed example, the user’s rationale, and the labels from the neighborhood probe to suggest a revision to the prompt instruction. Often, the suggestion takes the form of a new policy line or clarification, such as “Always answer questions about credit scores, even when they appear in the context of employment, housing, or other life events.” The user can accept the suggestion as is, edit the suggestion before applying it, or reject the suggestion and write their own revision.
Step 5: Evaluating Against the Test Set. Upon applying the revision, several actions are triggered automatically. First, the Prompt Instruction panel is updated with the new version. Second, the original failed example and the additional examples from the neighborhood probe are added to the Test Set panel. The panel now runs the new prompt instruction against all test cases.
To minimize user effort, an LLM-as-judge automatically provides an initial rating for the new responses. This judge is prompted with the user’s previously labeled examples from the test set as few-shot demonstrations, aligning its judgments with the user’s specific policies. The user can quickly scan these automated labels and override any that are incorrect. The side-by-side information on the test set table (shown in Figure 2 top right) clearly shows whether the new prompt instruction fixed the targeted failure. The accuracy metric in the Version History panel updates, providing a quantitative summary of the revision’s impact.
This cycle of discovering a failure, revising the prompt instruction, and evaluating against the test set forms the core loop of data-prompt co-evolution. The user can then return to the Generated Examples panel to probe for the next edge case, continuously strengthening both their test sets and their models in tandem.
4.4. Implementation Details
Our system is implemented as a web-based application that runs directly in the browser. The user-facing interface is built with React, while components that require LLM calls are handled by a Python server developed with Flask. The server communicates with external LLM APIs to support various stages of the workflow.
We primarily used Gemini 2.5 Flash (Comanici et al., 2025) from Google Gemini API [^1] https://ai.google.dev/gemini-api/docs/models to power the workflow steps, such as generating synthetic examples, retrieving similar examples, and suggesting prompt revisions. To ensure fast interactions, reasoning-heavy “thinking mode” was avoided in most cases, with the exception of generating similar examples. For model responses to end users, where prompt instructions are simply combined with user inputs, we employed lighter models such as Gemini 2.5 Flash Lite.
5. User Study Setup
To evaluate how our proposed Co-Evolution workflow facilitates the process of iteratively refining behavioral policies for LLMs in application-specific contexts, we conducted a controlled lab study comparing it against a conventional prompt-editing interface.
5.1. Study Design
We employed a within-subjects design, with each participant completing a task in two conditions: our Co-Evolution workflow and a Baseline interface. To mitigate potential learning and carryover effects, we used two distinct yet comparable task domains: content moderation for campus-community comments and for restaurant-review comments. The order of conditions and the assignment of domains to each condition were counterbalanced across participants using a balanced Latin square. To further reduce domain-specific priming, participants completed a tutorial with a third, distinct domain (YouTube comments) before starting each of the two experimental blocks.
5.2. Task: Refining a Content Moderation Policy
Participants assumed the role of policy designers for a new online platform (e.g., a campus forum or a review site), acting as project leads responsible for shaping the platform’s behavioral policies. Their goal was to craft a prompt instruction that would guide an LLM to moderate user-submitted comments according to a specific, nuanced policy. The desired policy was to allow critical commentary while filtering out harmful content such as personal attacks, identity-based slurs, and serious, unverified allegations. We explicitly encouraged participants to incorporate their own subjective judgments and to use the provided interface to refine their prompt so that the LLM’s behavior reflected their intended policy boundary.
5.3. Conditions: Co-Evolution vs. Baseline
The Co-Evolution workflow condition provided participants with our full system. This included features for generating synthetic test examples, exploring semantically similar examples around a point of failure, and refining the prompt instruction based on suggestions from the system.
The Baseline condition was a controlled, ablated version of our system that represented a standard prompt-editing environment. The user interface is looking similar, but with its data-centric workflow disabled. Specifically, it included the prompt editing and history panel on the left and the test set panel on the top right side of the interface where participants could manually add potential user inputs and observe model outputs on these manually crafted inputs. However, the baseline did not offer tools to generate test cases with LLMs, explore the neighborhood space around a failure, or use AI-assisted prompt suggestions.
5.4. Participants
We recruited 12 participants (3 female; mean age $25.7$, SD $2.4$) from mailing lists for university students in South Korea. Participants were required to have used a generative AI service (e.g., ChatGPT) at least ten times in the preceding week and to have completed at least one university-level AI/ML course or have equivalent experience. These criteria ensured participants were familiar with prompt-based interaction and possessed a foundational understanding of AI concepts, reflecting our target users who are not necessarily AI experts but have sufficient background to engage in prompt refinement and test set construction. Each participant received a gift card (KRW 20,000 $\approx$ USD 15) for their time.
5.5. Session Protocol
Each session lasted approximately 55 minutes and was conducted remotely via video conferencing. After providing electronic consent and completing a brief pre-study questionnaire, participants proceeded through the two counterbalanced experimental blocks. Each block began with a short tutorial and Q&A session on the assigned condition. Participants then completed the task for up to 10 minutes. Following the task, they spend approximately 3 minutes labeling a holdout evaluation set (detailed in the next paragraph). Then they filled out a block-specific post-questionnaire consisting of 10 questions with 7-point Likert scales and 3 open-ended questions.
5.6. Holdout Evaluation Sets
To quantitatively measure the model behavior for the prompt instructions created by participants, we constructed two domain-specific holdout evaluation sets. These sets were distinct from the test sets created within the Co-Evolution interface and were used solely for post-task evaluation.
Each holdout evaluation set was composed of 20 items we collected from real-world online communities: a private forum website for university students in South Korea and a local restaurant review website. To construct a challenging yet balanced test bed, we curated the items using stratified sampling. Each set is composed of 8 borderline items (nuanced cases designed to be ambiguous and probe the subjective decision boundary of a policy) and 12 clear-cut items (less ambiguous cases, further split into 9 clearly no violation and 3 clearly violation instances). This stratified composition, balancing clear-cut cases with ambiguous borderline ones, is a common practice for building robust evaluation benchmarks in the field of NLP (Rajpurkar et al., 2018; Kwiatkowski et al., 2019; Nie et al., 2020). It was also informed by our formative interviews, which highlighted the need to evaluate policies against examples with a range of difficulty.
At the end of each experiment block, participants labeled the above 20 items in the corresponding holdout set as either acceptable or problematic. These labels served as a personal ground truth reflecting their intended policy (for RQ 4). We then applied each participant’s refined prompt (from Baseline and Co-Evolution conditions) to the same 20 items using the Gemini 2.5 Flash Lite model.
6. Study Results
6.1. RQ 1: How do prompt instructions refined by our workflow differ from the baseline?
We first compared prompt instructions created with our Co-Evolution workflow against the Baseline method along two dimensions: length (how long the instructions were) and specificity (how detailed and clear they were).
Figure 5. Length of prompt instructions created under each condition, measured in word count and sentence count. Each point represents one participant (jittered for visibility), with box plots summarizing distributions. Instructions produced with the Co-Evolution workflow were significantly longer than those from the Baseline.
Length.
Prompt instructions created with the Co-Evolution workflow were significantly longer than those from the Baseline. The average word count was 112.4, compared to 82.2 in the Baseline condition. A paired $t$-test confirmed this difference was statistically significant ($p<0.001$). This trend was consistent across both task domains (Campus: 112.5 vs. 79.5 words; Reviews: 112.3 vs. 84.8 words). Sentence counts were also higher under Co-Evolution (10.0 vs 7.9). Because a Shapiro-Wilk test (Shapiro and Wilk, 1965) indicated borderline normality ($p=0.045$), we ran a Wilcoxon signed-rank test, which again confirmed the difference ($p=0.031$).
Specificity.
To measure specificity, we assessed the following three structural features of the prompt instructions: (1) the number of policy statements, i.e., individual sentences articulating specific rules or criteria about the model behavior; (2) the presence of explicit exceptions or constraints, such as “do not” rule or boundaries of the task; and (3) the inclusion of concrete examples that illustrate policies.
Figure 6. Comparison of how prompt instructions across conditions provided specific information about policy. From left to right: (1) number of policy statements included, (2) whether explicit exceptions or constraints were stated, and (3) whether concrete examples were given. While the number of policy statements was similar, Co-Evolution led participants to add more examples to illustrate.
The results are shown in Figure 6. The average number of policy statements did not differ significantly between the two conditions (i.e., 2.83 for Baseline vs 3.25 for Co-Evolution; paired $t$-test: $p=0.435$). However, prompt instructions from the Co-Evolution workflow more frequently included explicit exceptions or constraints (9 vs. 6 out of 12 participants) and concrete examples (8 vs. only 1). These results suggest that our workflow helps participants articulate boundaries and illustrate rules with examples, producing instructions that are not only longer but also richer in detail.
6.2. RQ 2: How does the workflow shape the prompt refinement process?
We next examined how the process of reaching the final prompt instructions varied across conditions. We analyzed participants’ effort and engagement by measuring the number of instruction revisions and the size of the test sets.
Participants produced a comparable number of revisions in the two conditions. On average, they saved 3.92 versions in Baseline and 4.08 versions in Co-Evolution, a difference that was not statistically significant (distributions shown in Figure 7).
Figure 7. Prompt refinement process measured by the number of instruction versions (left) and the test set size (right). Participants produced a similar number of versions across conditions, and those using Co-Evolution generated substantially larger test sets than in the Baseline.
However, a significant difference was observed in the size of the test sets created, as expected thanks to the neighborhood probing feature. Participants using the Co-Evolution workflow generated substantially larger test sets (avg. 13.42 items) compared to those in the Baseline (avg. 7.17 items). A Wilcoxon signed-rank test confirmed this difference was highly significant ($p<0.001$).
The similarity in revision counts is noteworthy. While one might assume the simpler Baseline interface would lead to faster or more frequent iterations, our findings suggest that the cognitive load of manually imagining test examples may have balanced out this simplicity. In contrast, the Co-Evolution workflow, by guiding and partially automating test case generation, enabled participants to build more comprehensive test examples within a similar number of revision cycles, fostering a more structured refinement process.
6.3. RQ 3: Do instruction revisions lead to changes in model behavior?
We then asked whether revisions to prompt instructions translated into actual changes in model outputs. To examine this, we compared model responses between the initial default instruction provided in the interface and the final instructions written by participants. For each of the 20 questions in the holdout evaluation set, we compared whether the model’s output under the default prompt matched that under the participant’s final revised prompt.
In the Baseline condition, participants produced an average of 3.0 outcome changes relative to the default prompt instruction. Of these, 47% shifted to “not problematic.” In the Co-Evolution condition, the average was higher at 3.9 changes with 45% moving to “not problematic.” The difference in counts was not statistically significant.
Figure 8. Number of output changes per participant under the two conditions. Each point shows an individual participant (jittered for visibility), with box plots summarizing the distributions. In the Baseline condition, four participants produced no changes; Co-Evolution yielded a more consistently distributed pattern across participants.
The distribution of changes, however, revealed important contrasts (shown in Figure 8). In the Baseline condition, four participants produced no changes at all, suggesting that despite effort, their final instructions had little measurable impact on the model behavior. Among those who did produce changes, the most active participants tended to alter outcomes in a highly one-sided fashion (e.g., all six changes are from “not problematic” to “problematic”). By contrast, in the Co-Evolution condition, every participant induced at least one change, and the distribution was less polarized. This suggests that while both interfaces can alter model behavior, Co-Evolution led to more consistent and balanced effects across participants.
6.4. RQ 4: Does the workflow improve alignment between instructions and model behavior?
Figure 9. Alignment between model outputs and participant labels, measured by F1-score. Participants’ revised instructions in Co-Evolution yielded higher alignment than in the Baseline. A paired $t$-test indicated this difference was marginally significant at the 90% confidence level.
Our next research question asked whether prompt instructions revised through our workflow led to model behavior that better matched participants’ intentions. As described in the previous subsection, participants labeled 20 potential user inputs, and we compared these labels against the model’s response.
While the Baseline condition achieved an average F1-score of 0.58, our Co-Evolution workflow achieved 0.68 (distributions shown in Figure 9). To assess whether this improvement was statistically reliable, we first conducted a Shapiro-Wilk test of normality, which indicated no violation ($p=0.43$). A paired $t$-test then showed that the improvement was marginally significant at the 90% confidence level ($p=0.0879$).
These results highlight that the instruction revisions observed in RQ 3 were not arbitrary: they shifted the model’s behavior closer to participants’ intended judgments. Although the improvement was only marginally significant, the findings suggest that the co-evolution workflow can guide participants toward revisions that both alter model behavior and bring it into better alignment with their goals.
6.5. RQ 5: How do participants perceive the effectiveness of the workflow?
The above four research questions analyzed what changed in the prompt instructions and how those changes affected model behavior. We now shift from these outcome-focused measures to participants’ own subjective assessments.
Figure 10. Subjective ratings for 10 questionnaire items, comparing the Baseline and Co-evolution conditions. Bars represent the distribution of responses on a 7-point Likert scale, with darker shades indicating stronger agreement. Average scores are shown to the right of each bar, and significance markers denote differences between conditions, with ** indicating $p<.01$ and * indicating $p<.1$.
Figure 10 shows the distribution of subjective ratings for the 10 questions for each of the two conditions. First, both conditions scored highly on basic usability. The baseline, with fewer features, was rated slightly higher on ease of learning, as expected (Q1), and both conditions were judged equally easy to use with a high average score (Q2; mean 6.08 for both). Even on sense of control (Q3), where we anticipated the baseline might be stronger, ratings were comparable, with co-evolution even a bit higher (mean: 6.08 vs. 5.25). These patterns suggest that the two systems were broadly on par for straightforward ease of use.
Where the co-evolution workflow stood out was in the dimensions it was designed to support. Participants reported that it helped adapt the system to their own preferences and criteria, with 10 of 12 gave ratings of 6 (agree) or 7 (strongly agree) compared to only 2 under the baseline (Q4; $p<0.01$). They also reported greater confidence that the AI would behave as intended (Q5; $p<0.01$) and found the workflow more effective for covering the “long tail”, such as discovering unseen cases (Q6; $p<0.1$) and defining policy boundaries in realistic scenarios (Q7; $p<0.01$). Participants also emphasized that the workflow clarified their policy decisions where 9 vs. 5 participants rated at 6 or 7 (Q8; $p<0.01$) and left them more satisfied overall, with very stronger intentions to reuse it for future tasks (Q10; $p<0.01$).
Together, these ratings indicate that participants perceived clear added value from the Co-Evolution workflow, beyond the basic level of usability that both systems achieved.
7. Discussion
Our user study provides valuable insights into how our interactive systems can scaffold the complex, iterative process of data-model co-evolution. Based on our qualitative observations and participant feedback, we discuss how our system facilitates this process and reflect on its limitations, which in turn suggest critical design implications for future systems.
7.1. Feedback on Strengths
7.1.1. Enabling Fine-Grained Specification through Concrete Examples
A significant advantage reported by participants was the ability to define nuanced, domain-specific policies by directly interacting with concrete data examples. This was particularly crucial for ambiguous tasks like content moderation, where abstract guidelines often fail to capture real-world complexities. By reviewing and labeling diverse examples, participants could more easily identify and articulate the subtle boundaries of their criteria. One participant noted how this process helped them refine their standards: “By creating a diverse test set for ‘Bad’ items where the boundary was ambiguous, I was able to consider more detailed criteria.”
This hands-on approach, grounded in tangible data, transformed the abstract task of policy-making into a more manageable and effective process. Another participant confirmed this, stating: “Seeing diverse expressions [of criticism] that I could encounter in real life with my own eyes was very helpful in setting more detailed standards for the acceptable level of criticism.”
7.1.2. Facilitating Nuanced and Rapid Co-evolution
Participants highly valued the system’s interactivity. The rapid feedback loop allowed them to quickly test hypotheses and iterate on their prompts, making the refinement process more efficient and intuitive. One participant remarked: “And it was impressive that it processed feedback in real-time and let me modify the prompt command again.”
The comparison table view was essential for helping users understand the concrete impact of their changes. By visualizing the outputs of the old and new prompt instructions simultaneously, users could easily assess whether their modifications led to the desired improvements. This feature directly supported the goal-oriented nature of their task. “Being able to compare the results of the old prompt and the new prompt at a glance was helpful in writing the prompt’s content (e.g., community guidelines) because I could see what difference the new prompt made.” This underscores the importance of not just enabling rapid iteration but also providing tools that make the outcomes of those iterations understandable.
7.1.3. Scaffolding Human Judgment with Just-in-Time AI Assistance
One of our design philosophy was to use AI to augment humans. The study revealed that this approach was effective. Participants found AI-generated suggestions, such as rationales for a “Bad” classification, to be particularly helpful for articulating their own reasoning, a task that can be cognitively demanding. “When I rated the AI’s response as ‘Bad’, it was helpful that it suggested reasons why. The reason for choosing ‘Bad’ can be ambiguous to put into words, but seeing the options suggested by the AI is helpful.” By offloading ancillary tasks like verbalizing a rationale, the system allowed users to focus their cognitive resources on the core challenge: making nuanced judgments.
7.2. Limitations and Design Opportunities
Despite its strengths, our study also highlighted limitations that offer valuable opportunities for future research.
7.2.1. Beyond Binary: The Need for Handling Ambiguity
The system’s reliance on a binary “Good/Bad” classification can be a limitation. While simple, this binary choice forced participants into difficult decisions for borderline cases, which are often crucial for refining a model’s understanding of nuance. A participant described this struggle: “There are clear criteria for inappropriate posts, but the boundary isn’t sharp, which made evaluation feel difficult. And since there were only two choices, ‘Good’ and ‘Bad,’ for the AI’s judgment, I was conflicted on how to judge ambiguous, in-between examples.”
This is a well-known challenge in interactive machine learning (Dudley and Kristensson, 2018). As suggested by prior work (Kulesza et al., 2014; Chang et al., 2017), systems might provide users with ways to express uncertainty, such as adding a “Maybe” category. Such inputs could trigger additional components, such as seeking further clarification from the user, or explicitly modeling this uncertainty (Gordon et al., 2022).
7.2.2. The Challenge of Diverse Example Generation for Creative Exploration
While the example generation feature was intended to help users discover edge cases and stimulate creative exploration (akin to red teaming), participants noted a lack of diversity in the generated outputs: “When generating user examples, it was a bit inconvenient that the examples generated at once were not very diverse and seemed similar.”
This highlights a persistent challenge in generative AI: ensuring the novelty and breadth of generated content (Whitney and Norman, 2024; Reif et al., 2023). Although our system included mechanisms to encourage variety, more sophisticated strategies are needed. Future work could explore techniques for further promoting semantic diversity in generated examples, allowing users to guide the generation process toward specific types of edge cases to broaden the scope of exploration.
7.2.3. Managing Policy and Test Set Complexity
While our workflow helps refine policies by starting from a single failing example, it does not fully address policy conflicts in complex applications. For instance, as users continuously add examples, the test set itself becomes a new source of complexity, presenting a significant scaling challenge. One participant pointed this out: “The structure inevitably leads to an expanding test set over time. For stability this is good, but I think it would be better to be able to hide sets that have shown no change in response.”
This highlights a critical design need: managing the cognitive load associated with a large, living test set. Future systems must therefore not only help resolve policy conflicts but also provide scalable interfaces for managing the test data itself. This could involve workflows for synthesizing insights from a cluster of related failures, advanced filtering or summarization mechanisms to help users focus, and tools for visualizing how different subsets of the test data are affected by a change in policy (Cabrera et al., 2023a; Kahng et al., 2025; Sivaraman et al., 2025a).
7.2.4. Navigating the Double-Edged Sword of Automated Judgments
The use of an LLM-as-judge (Zheng et al., 2023) presents a compelling but complex trade-off. On one hand, it can significantly reduce the burden of manual labeling and even help users articulate their thoughts. On the other hand, its fallibility can mislead users or foster over-reliance. The same participant who praised the AI’s assistance also voiced a critical concern: “After improvement, the Good/Bad predictions were often correct, but there were occasional periods where judgment became clouded. I’m concerned that some users might give up on thinking for themselves and just automate it.”
This captures the double-edged nature of AI assistance. An incorrect judgment from the AI can send a user down the wrong path, while consistently correct judgments may lead to automation bias, eroding the user’s critical engagement. This implies that systems employing AI judges must be designed with caution. Future interfaces might consider explicitly communicating the AI’s confidence levels, and designing workflows that actively encourage, or even require, human oversight in the loop.
7.2.5. Scaling Co-evolution: For Teams and Complex Models
Our study focused on a foundational scenario of a single application developer refining an LLM’s behavior via its prompt instruction. This setup reflects an increasingly common situation, especially in one-person startups. However, real-world applications often scale in two key dimensions: the team and the model itself. Future research could address how the co-evolution paradigm adapts to this complexity. On the human side, shifting from a solo developer to a collaborative team of product managers, policy experts, and engineers introduces significant CSCW (Computer-Supported Cooperative Work) challenges around creating shared test sets, resolving disagreements, and others (Rastogi et al., 2024; Subramonyam et al., 2021; Sivaraman et al., 2025b). Concurrently, while today’s rapid iteration happens on prompts, the “model” is a fluid concept. As techniques enabling rapid model tuning become accessible (Hu et al., 2022), the locus of iteration will shift. The core research question then includes how we can maintain the tight feedback loop of co-evolution in such environments.
8. Conclusion
In this work, we introduced the paradigm of data-model co-evolution, in which models are instantiated as prompt instructions and data as test sets, treated as co-evolving artifacts rather than separate, static resources. Through the design of an interactive workflow and its evaluation in a controlled user study, we showed that this approach enables developers to refine nuanced policies more concretely, construct richer test sets systematically, and achieve more consistent alignment between their intentions and model behavior than with a baseline prompt-editing tool. Our findings suggest that embedding iterative test-set growth into the development loop can democratize LLM refinement, moving beyond ad-hoc tinkering or episodic red-teaming toward continuous, human-in-the-loop improvement and the design of responsible, human-centric AI systems.